OpenStudy (anonymous):

A student finds the density of a liquid by measuring its mass and its volume. The following is a summary of his measurements. mass of empty beaker = (20 ± 1) g mass of beaker + liquid = (70 ± 1) g volume of liquid = (10.0 ± 0.6) cm3 He correctly calculates the density of the liquid as 5.0 g cm–3. What is the uncertainty in this value? A 0.3 g cm–3 B 0.5 g cm–3 C 0.6 g cm–3 D 2.6 g cm–3

OpenStudy (anonymous):

Is answer B

OpenStudy (anonymous):

(2/50+0.6/10)*5=0.5

OpenStudy (anonymous):

density= mass/volume when we are dividing or multiplying we add % uncertainties so, % uncertainty in mass= ( 2*1/70-20)*100= 4% ...................... volume= (.6/10)*100= 6% add these two and u get 10% Multiply it with the value of density and u get 0.5 :)